40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 180 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  8921   Thu Jul 25 02:53:00 2013 KojiUpdateGeneralIFO Alignment after TT flipping - no progress

There was no progress tonight after Jenne left.
I could not find any reasonable fringes of the IFO after 3 hours of optics jiggling.

* I jiggled TT1 and TT2. The slider has not been restored.
We should probably look at the value in the day time and revert them.
(Still this does not ensure the recovery of the previous pointing because of the hysteresis)

* The arms are still aligned for the green.
It's not TEM00 any more because of the vent/drift but the fringe is visible (i.e. eigenaxis is on the mirror)

* As we touched PR3, the input pointing is totally misaligned.

To Do / Plan

* We need to find the resonance of the yarm by the input TTs. Once the resonance is found, we will align the PRM.

* Move the BS to find the xarm resonance.

* Finally align SRM

* It was not possible to find the resonance of the yarm without going into the chamber. Definitely we can find the spot on the ITMY by a card, but we are not sure the beam can hit the ETMY. And the baffles makes the work difficult.

* One possibility is to align the input beam so that the ITMY beam is retroreflected to the PRM. I tried it but the beam was not visible form the camera.

  8920   Wed Jul 24 22:58:03 2013 CharlesUpdateISSISS - Full Schematic - Updated

 I have made significant changes to the ISS schematic, mostly in the form of adding necessary subsystems.

Some changes I have made:

  • Added a front page with sheet symbols that are representations of the other schematic sheets.
  • Added an 'Excitation' subsystem for use in determining the closed-loop transfer function
  • Added an instrumentation amplifier (with ADA4004s at Rana's recent suggestion) to handle the differential input from the PD
  • Included a switchable inverting amplifier (Gain of 1 or -1) to ensure we have the correct polarity
  • Made it so the first filtering stage is immediately active when the ISS loop is closed
  • Added LP filters with large time constants to buffer/delay trigger signals
  • Added test points all over the board
  • Refined a few buffer amplifiers

On the front page, all inputs and outputs are currently BNC ports, although this is most likely not the final design that will be used. For instance, the ports ENABLE, INPUT GND and INVERT are supposed to be logic inputs for a MAX333a switch. These will most likely be front panel switches that either connect the switch's logic pin to GND (Logic 0) or something like a +5 V supply (Logic 1).

I also have not included power regulation for my board although I have some of the actual D1000217 Chasis Power Regulator boards and I'll incorporate those in my design soon.

Attachment 1: 40mServo_v1.pdf
40mServo_v1.pdf 40mServo_v1.pdf 40mServo_v1.pdf 40mServo_v1.pdf 40mServo_v1.pdf
  8919   Wed Jul 24 19:21:56 2013 JamieHowToSUSSUS MEDM screen modernization

I started poking around at what we want for new SUS MEDM screens.  Rana and I decided we'd start with the ASC TIPTILT screens:

newsusmedm.png

It's missing some things (like SIDE OSEMS) but it should provide a good starting point.

I copied the entire <userapps>/asc/common/medm/asctt directory to a new directory in our sus area:

controls@rossa:/opt/rtcds/userapps/release 0$ cp -a asc/common/medm/asctt sus/c1/medm/new

I then removed all the useless file name prefixes.  We still need to go through and sed out all the ASC stuff in the MEDM files themselves.

It makes heavy use of macro substitution, which is good (it's what we're using now).  So once we clean up all the channel names, we should just be able to swap out the pointers in our overview screens to the new screens (or rename things).  In the mean time, during development, you can run:

controls@rossa:/opt/rtcds/userapps/release 0$ medm -x -macro "IFO=C1,ifo=c1,OPTIC=ITMX" sus/c1/medm/new/OVERVIEW.adl 

  8918   Wed Jul 24 15:07:54 2013 KojiUpdateSUSSR2 flipped

After the first flipping, X/Y arms were aligned and locked. Then the ASS aligned the arms.

  8917   Wed Jul 24 14:26:24 2013 ranaUpdateIOOFound the cause of mysterious MC motion

Yes, this was not ELOG'd by me, unfortunately. This was the MC tickler which I described to some people in the control room when I turned it on.

As Koji points out, with the MCL path turned off this injects frequency noise and pointing fluctuations into the MC. With the MCL path back on it would have very small effect. After the pumpdown we can turn it back on and have it disabled after lock is acquired. Unfortunately, our LOCKIN modules don't have a ramp available for the excitation and so this will produce some transients (or perhaps we can ezcastep it for now). Eventually, we will modify this CDS part so that we can ramp the sine wave.

  8916   Wed Jul 24 13:41:13 2013 JenneUpdateSUSSR2 flipped

[Jenne, Annalisa]

SR2 is flipped, and reinstalled.  We did that before lunch, and we're about to go in and work on SR3 and PR3.

EDITS / Notes:

I set dog clamps to have a reference position of where the tip tilt was, then I removed SR3 from the chamber.  Once out, I followed the same procedure I used for PR2 during the last vent - I removed the whole suspension (top mount, wires, optic) from the cage, and laid it down flat.  Then I loosened the set screw which pushes on the teflon nudge, removed the mirror, inspected it, and put it back in, with the HR side facing the back side of the ring.  Then I replaced the suspension system in the cage, and put the mirror back into the chamber. 

When I loosened the teflon nudge at the top of the mirror holder ring, the optic seemed to fall down a tiny bit.  I think this implies that the HR surface of the optic did not used to be parallel to the front face of the mirror holder ring.  When I put the suspension back onto the cage, the pitch balancing was very bad.  We checked the level of the table that I had the cage on, and it was miraculously pretty level, so I did the pitch balancing out of the chamber. 

Also, during my quick inspection of the mirror (not thorough, just using room lights), I noticed a small fleck of lint near the edge of the optic on the HR surface.  The HR surface is now on the outside of the SRC, but we should still blow at the optic with the ionized nitrogen to get it off.

I did not think to check the fine-tuning alignment of SR2....Koji did that after lunch (which I will elog about in a separate elog).

 

  8915   Wed Jul 24 10:35:41 2013 SteveUpdateVACBS, ITMY doors are removed

Quote:

We will open the BS and ITMY doors first thing tomorrow morning. I plan to try to be in around 9 am. The first order of business will be to flip the folding mirrors that are not currently flipped (SR2, SR3, PR3).

 Jenne, Annalisa & Steve

Attachment 1: beforeDoorsOff.png
beforeDoorsOff.png
Attachment 2: particlecount10d.png
particlecount10d.png
  8914   Tue Jul 23 22:55:13 2013 JenneUpdateVACBS, ITMY doors to be opened in the morning

We will open the BS and ITMY doors first thing tomorrow morning. I plan to try to be in around 9 am. The first order of business will be to flip the folding mirrors that are not currently flipped (SR2, SR3, PR3).

  8913   Tue Jul 23 21:32:43 2013 KojiUpdateIOOFound the cause of mysterious MC motion

Thesedays we were continuously annoyed by unELOGGED activities of the interferometer.

MC2 LOCKIN was left on and has continuously injected frequency noise and beam pointing modulation
during all of the comissioning / vent preparation.

C1:SUS-MC2_LOCKIN2_OSC_FREQ was 0.075
C1:SUS-MC2_LOCKIN2_OSC_CLKGAIN was 99

For more than a week ago we noticed that the curve of the MC WFS stripchart suddenly got THICKER.
MC WFS, arm transmission, beam pointing... everything was modulated.
It was not WFS instability, and it was not the cavity mirrors.

Today I made the investigation and finally tracked down the cause of this issue to be on MC2 suspension.
Then it was found that this LOCKIN was ON.

There is no direct record of this lockin in the frame files.
From the recorded channel "C1:IOO-WFS2-YAW_OUT16" (which is the trace on the StripTool chart on the wall)
It was turned on at July 10th, 2:00UTC (July 9th, 7PM PDT)

  8912   Tue Jul 23 20:41:40 2013 gautamConfigurationendtable upgradeFull range calibration and installation of PZT-mounted mirrors

 Given that the green beam is to be used as the reference during the vent, it was decided to first test the PZT mounted mirrors at the X-endtable rather than the Y-endtable as originally planned. Yesterday, I prepared a second PZT mounted mirror, completed the full range calibration, and with Manasa, installed the mirrors on the X-endtable as mentioned in this elog. The calibration constants have been determined to be (see attached plots for aproximate range of actuation):

M1-pitch: 0.1106 mrad/V

M1-yaw: 0.143 mrad/V

M2-pitch: 0.197 mrad/V

M2-yaw: 0.27 mrad/V


Second 2-inch mirror glued to tip-tilt and mounted:

  • The spot sizes on the steering mirrors at the X-end are fairly large, and so two 2-inch steering mirrors were required.
  • The mirrors already glued to the PZTs were a CVI 2-inch and a Laseroptik 1-inch mirror.
  • I prepared another Laseroptik 2-inch mirror (45 degree with HR and AR coatings for 532 nm) and glued it to a PZT mounted in a modified mount as before.
  • Another important point regarding mounting the PZTs: there are two perforated rings (see attached picture) that run around the PZT about 1cm below the surface on which the mirror is to be glued. The PZT has to be pushed in through the mount till these are clear of the mount, or the actuation will not be as desired. In the first CVI 2-inch mirror, this was not the the case, which probably explains the unexpectedly large pitch-yaw coupling that was observed during the calibration [Thanks Manasa for pointing this out]. 

Full range calibration of PZT:

Having prepared the two steering mirrors, I calibrated them for the full range of input voltages, to get a rough idea of whether the tilt varied linearly and also the range of actuation. 

Methodology:

  • The QPD setup described in my previous elogs was used for this calibration. 
  • The linear range of the QPD was gauged to be while the output voltage lay between -0.5V and 0.5V. The calibration constants are as determined during the QPD calibration, details of which are here.
  • In order to keep the spot always in the linear range of the QPD, I stared with an input signal of -10V or +10V (ie. one extreme), and moved both the X and Y micrometers on the translational stage till both these coordinates were at one end of the linear range (i.e -0.5V or 0.5V). I then increased the input voltage in steps of ~1V through the full range from -10V to +10V DC. The signal was applied using a SR function generator with the signal amplitude kept to 0, and a DC offset in the range -5V to 5V DC, which gave the desired input voltages to the PZT driver board (between -10V DC and 10V DC).
  • When the output of the QPD amp reached the end of the linear regime (i.e 0.5V or -0.5V), I moved the appropriate micrometer dial on the translational stage to take it to the other end of the linear range, before continuing with the measurements. The distance moved was noted. 
  • Both the X and Y coordinates were noted in order to investigate pitch-yaw coupling.

Analysis and remarks:

  • The results of the calibration are presented in the plots below. 
  • Though the measurement technique was crude (and maybe flawed because of a possible z-displacement while moving the translational stage), the calibration was meant to be rough, and I think the results obtained are satisfactory. 
  • Fitting the data linearly is only an approximation, as there is evidence of hysteresis. Also, PZTs appear to have some drift, though I have not been able to quantify this (I did observe that the output of the QPD amp shifted by an amount equal to ~0.05mm while I left the setup standing for an hour or so).  
  • The range of actuation seems to be different for the two PZTs, and also for each degree of freedom, though the measured data is consistent with the minimum range given in the datasheet (3.5 mrad for input voltages in the range -20V to 120V DC). 

 

PZT Calibration Plots

The circles are datapoints for the degree of freedom to which the input is applied, while the 'x's are for the other degree of freedom. Different colours correspond to data measured with the position of the translational stage at some value.

                                            M1 Pitch                                                                                             M1 Yaw

M1_Pitch_calib.pdf     M1_Yaw_calib.pdf

 

                                              M2 Pitch                                                                                        M2 Yaw 

M2_Pitch_calib.pdf     M2_Yaw_calib.pdf

 



Installation of the mirrors at the X-endtable:

The calibrated mirrors were taken to the X-endtable for installation. The steering mirrors in place were swapped out for the PZT mounted pair. Manasa managed (after considerable tweaking) to mode-match the green beam to the cavity with the new steering mirror configuration. In order to fine tune the alignment, Koji moved ITMx and ETMx in pitch and yaw so as to maximise green TRX. We then got an idea of which way the input pointing had to be moved in order to maximise the green transmission.

 

Attachment 5: PI_S330.20L.pdf
PI_S330.20L.pdf
  8911   Tue Jul 23 19:38:58 2013 gautamUpdateCDSCharacterisation of DAC at 1X9

 

 I just finished carrying out the same checks for the DAC at 1X9 (with channels 9 through 16 that are unused as of now) as those I had done for the DAC at 1Y4, as the hardware prep up till now was done with the characterisation of the DAC at 1Y4. Conclusions:

  • The accessible range of output voltage are -10 V to +10V w.r.t ground --> No change needs to be made to the gain of the HV amplifier stage on the PZT Driver Board
  • The pin-outs of the DAC Adaptor Board at 1X9 is identical to that at 1Y4 --> Custom ribbons do not need to be modified.
  • The PSD of the DAC output has a peak at 64 kHz --> Notches on AI Board do not need to be moved again.

I will now proceed to install various pieces of hardware (AI Board, PZT driver board, HV Power Supply and cabling) at 1X9, while not making the connection to the PZTs till I receive the go ahead. 

  8910   Tue Jul 23 19:37:08 2013 AnnalisaUpdateendtable upgradeEnd table picture

Quote:

The spot on the IPANG QPD was checked. The spot is higher than the center and South side of the lens.
Some photos are found below.

The spot on the IPANG steering mirrors in the ETMY chamber was also checked.
It is clipped at the top of the steering mirror. (See attachment 4)
So basically the spot is about 1" above the center of the mirror.

 After the vent, the IPang spot position on the steering mirrors on the Yend table moved approximately by 1 inch down.

Inside the chamber, the spot position is in the center of the steering mirror. (difficult to take a picture because the PSL beam power has been reduced)

Attachment 1: IMG950073.jpg
IMG950073.jpg
Attachment 2: IMG950072.jpg
IMG950072.jpg
  8909   Tue Jul 23 16:47:01 2013 gautamUpdateCDSExcitation points set up on c1scx

 I just compiled and installed the model with the excitation points on c1scx and then restarted framebuilder. The channels I set up are now showing up in the awggui dropdown menu. I will do the tests on the DAC channels shortly.

Just to keep things on record, these are the steps I followed:

  • opened the model c1scx (path: /opt/rtcds/userapps/release/sus/c1/models) with MATLAB
  • Added 8 excitation points and saved the model. A copy has been saved as c1scx.mdl.r2010b because of the recent upgrade to r2013a. 
  • ssh to c1iscex (computer running the model c1scx). 
  • Entered the following sequence of commands in terminal: rtcds make c1scx ,  rtcds install c1scx , rtcds start c1scx 
  • ssh to framebuilder, and restarted the framebuilder by entering telnet fb 8088   and then   shutdown.
  8908   Tue Jul 23 16:39:31 2013 KojiUpdateGeneralFull IFO alignment recovered

[Annnalisa Koji]

Full alignment of the IFO was recovered. The arms were locked with the green beams first, and then locked with the IR.

In order to use the ASS with lower power, C1:LSC-OUTPUT_MTRX_9_6 and C1:LSC-OUTPUT_MTRX_10_7 were reduced to 0.05.
This compensates the gain imbalance between TRX/Y siganls and the A2L component in the arm feedback signals.

Despite the IFO was aligned, we don't touch the OPLEVs and green beams to the vented IFO.

Attachment 1: alignment.png
alignment.png
  8907   Tue Jul 23 14:00:03 2013 SteveUpdateVACthe vent is completed

 The 40m IFO has reached  atmospher in 5 hours. It is ready to open chamber condition. The RGA is pumped with the maglev.

P1 pirani gauge is contact dependent as you see it on the linear plot It will be replaced during this vent.

The venting speed was 2-4 Torr / min

Atm2 shows how the BS is sensing the venting air cylinder changes. 

The 4th cylinder of instrument  grade air  bump is overlapping with our janitor working at the BS chamber.

 

 

Attachment 1: Atm.png
Atm.png
Attachment 2: 5cyl.png
5cyl.png
Attachment 3: ventSum1d.png
ventSum1d.png
  8906   Tue Jul 23 13:55:08 2013 KojiUpdateIOOMC manually aligned

The MC was manually aligned. The spot positions were measured and it is consistent with the measurements done yesterday.

Attachment 1: MCalignment.png
MCalignment.png
Attachment 2: MCspot.png
MCspot.png
  8905   Tue Jul 23 13:29:58 2013 KojiSummaryGeneral 

record of the initial state

Attachment 1: initial_state_130723.png
initial_state_130723.png
  8904   Tue Jul 23 08:21:31 2013 SteveUpdateVACVent has started

 

Thanks for the good preparation.  IFO pressure P1=20 Torr

Attachment 1: beforVent.png
beforVent.png
Attachment 2: startVent.png
startVent.png
  8903   Tue Jul 23 04:31:36 2013 JenneUpdateGeneralVent preparation - Ready for Vent

After everyone's work today (good teamwork everybody!!), we are a GO for the vent.

Steve, please check the jam nuts, and begin the vent when you get in.  Thanks.

  8902   Tue Jul 23 04:26:54 2013 JenneUpdateGeneralMC spot positions

After Koji and I lowered the power into the PMC and saw that the MC locked nicely, I remeasured the spot positions (no alignment on the PSL table, or of the MC mirrors has been done.  Also, WFS are off, since there isn't any power going to them).

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[1.1999406656184595, 0.63492727550953243, 1.0769104750021909, -1.0260011922577466, -1.059439987970527, -1.2717741991488549]

MCdecenter_22July2013_lowpower.png

The spot positions seem to have actually gotten a bit better in pitch (although between 2 consecutive measurements there was ~0.5mm discrepancy), and no real change in yaw.  This means that Rana was right all along (surprise!), and that decreasing the power before the PMC reduces alignment pain significantly.

  8901   Tue Jul 23 04:22:03 2013 KojiUpdateGeneralVent preparation - In progress

[Koji Jenne]

Low power MC locking

- Rotated HWP right after the laser

- Put a knife edge beam dump at the output of the PBS after the HWP.

- Replaced the PO mirror for the MC refl by an HR mirror.

- PMC:
Input offset from 0 to 0.29
Servo Gain from 10 to 30
=> Transmission 0.84 (1.2W at the MC input) to 0.069 (100mW)

- MC:

VCO Gain from 25 to 31
MC REFL: Unlocked 3.6 Locked 0.38-0.40

  8900   Tue Jul 23 04:07:48 2013 gautamUpdateCDSExcitation points set up on c1scx

 

 In light of recent events and the decision to test the piezo tip-tilts for green beam steering on the X-end table, I have set up 8 excitation points to channels 8 through 15 of the DAC on c1scx (as was done earlier for the DAC at 1Y4 with Jenne's help) in order to verify that the pin-outs of the DAC interface board. I have not yet compiled the model or restarted the computer, and will do these tomorrow, after which I will do the test. The channels are named YYY_CHAN9 etc. 

 

 

 

 

  8899   Tue Jul 23 03:43:32 2013 manasaUpdateGeneralVent preparation - In progress

There seems to be an unexplained oscillation in X arm cavity transmission for IR when the cavity is locked using the POX error signal.

Their origin is not related to the oplevs because the oscillation does not exist when LSC is OFF and the arms are controlled only by the oplevs and OSEMs.

  8898   Tue Jul 23 03:10:57 2013 KojiUpdateGeneralVent preparation - In progress

Centering of the oplev beams: done

Recording the OSEM values: done

Attachment 1: OSEM_RECORD.png
OSEM_RECORD.png
Attachment 2: OPLEVS.png
OPLEVS.png
  8897   Tue Jul 23 01:30:27 2013 JenneUpdateASCPOP QPD analysis

I have some data for how much motion of any PRMI-relevant optic affects the beam seen by the POP QPD. 

For this, I am using the QPD calibration from the micrometer (elog 8851) to get me from counts to mm of motion.  Note that the pitch calibration hasn't been redone (I tried locking the PRMI this afternoon, but ITMX kept drifting away from me**, so I didn't get any more data.) The pitch calibration is obviously very rough, since I only have 2 points defining my fit line. 

Anyhow, if we assume that's close enough to get us started, I now have a calibrated QPD spectrum:

QPD_spectra_only.png

As detailed in elog 8854, I took single frequency transfer functions, to determine the effect at the QPD from shaking any single PRMI optic.  These transfer functions gave me a conversion factor between the optics' oplev readings (in microradians) to the counts seen at the QPD.  I used this number, as well as the QPD calibration from the micrometer data, to convert each optics' oplev spectra to motion that one would expect to see at the QPD. 

I have not yet completely figured out how to make an estimate of the PR folding optics' affect on the POP QPD spot position, if I know their motion.  The current plan is to do as Den did in elog 8451, and infer the PR2/3 motion from the ITMX/BS motion measured by the oplevs.  My plan was to take the spectra of the oplev signals while the BS/ITMX are undamped, divide by the SOS pendulum transfer functions, then multiply by the TT transfer functions (which I finally wrote down in elog 8564).  I'm planning on using the undamped data, since the oplev signals are still within the linear range of the oplev QPDs, and I won't have to take the SUS damping into account.  Anyhow, after I do that, I'll have an idea of how much the tip tilts are moving, but not what that does to the cavity axis.

However, after looking at the plots below, it seems like the PRM is the main culprit causing the PRC axis motion, although the BS (and to a smaller extent the ITMs) are not innocent.  Since the plots get very busy very quickly, I have many plots, each plot comparing one of the above QPD spectra (either pitch or yaw) with a single optics' oplev inferred motion.

EDIT:  After talking with Koji, I realize that, since the ASC was engaged during the PRM oplev spectrum measurement, I cannot yet say whether the motion is due to PRM, or if it is from PR2 or PR3, and imprinted on the PRM via the ASC servo.  The lump where the PRM-caused motion is greater than the QPD spectra is entirely in the region where the ASC is active.  So, the QPD motion I expect without the ASC would be something like the green trace in the PRM comparison plots.  The blue trace is then the closed loop measurement.  Since the ITMs and BS are below the closed loop values, they aren't the ones causing the big lump.  I should retake all of these spectra at a time when the PRMI is locked, but the ASC is not engaged.  I'm not sure if I'll have a chance to do that tonight or not.  If I can find some GPS times when the PRMI was locked, before we had ASC, I can get the oplev data.

PRM:

QPD_yaw_motion_from_PRM.pngQPD_pit_motion_from_PRM.png

BS:

QPD_yaw_motion_from_BS.pngQPD_pit_motion_from_BS.png

ITMX:

QPD_yaw_motion_from_ITMX.pngQPD_pit_motion_from_ITMX.png

ITMY:

QPD_yaw_motion_from_ITMY.pngQPD_pit_motion_from_ITMY.png

 I think part of the reason PRM is dominating is that it's damped motion is ~10x greater than any other optics', most noticeably the BS'.  I'll write a quick separate elog about this.  Also, note that the ~3Hz resonant gain had been turned off in the PRM oplev loop, but not in any other loops.  This is why there isn't the sharp dip in the PRM's oplev motion.  Also, since the PRM ASC was engaged for this measurement, and the ASC pushes on the PRM to minimize the QPD motion, it isn't totally crazy that the PRM's motion is greater than what we actually see at the QPD, if it is compensating for the motion of other optics.

 

** Re: PRMI locking this afternoon, it was almost as if ITMX were bi-stable.  I aligned both arms, to set the ITM positions.  Then, I would lock and tweak up the michelson to get the AS port nice and dark (usually touching ITMX today, since it seemed like the drifter....ITMX at this point was usually between -7 and -15 microradians in pitch from the center of the oplev QPD).  When I then brought the PRM back into alignment, ITMX was starting to drift away.  As soon as I hit the LSC Enable switch, and looked back over to the OpLev screen, ITMX was misaligned, usually around -65 urad in pitch.  I did this circus probably 3 or so times before giving up.  Koji said that he had seen this bi-stability before, but he didn't remember what fixed it.  The drifting that Koji mentioned in elog 8801 seems to have been fixed by centering all the PRMI oplevs every day, but I had already done that, and was still seeing ITMX drift.

  8896   Tue Jul 23 00:51:46 2013 KojiUpdateendtable upgradeEnd table picture

The spot on the IPANG QPD was checked. The spot is higher than the center and South side of the lens.
Some photos are found below.

The spot on the IPANG steering mirrors in the ETMY chamber was also checked.
It is clipped at the top of the steering mirror. (See attachment 4)
So basically the spot is about 1" above the center of the mirror.

Attachment 1: P7224222.JPG
P7224222.JPG
Attachment 2: P7224223.JPG
P7224223.JPG
Attachment 3: P7224224.JPG
P7224224.JPG
Attachment 4: P7224238.JPG
P7224238.JPG
  8895   Mon Jul 22 22:06:18 2013 KojiUpdateCDSFE Web view was fixed

FE Web view was broken for a long time. It was fixed now.

The problem was that path names were not fixed when we moved the models from the old local place to the SVN structure.

The auto updating script (/cvs/cds/rtcds/caltech/c1/scripts/AutoUpdate/update_webview.cron) is running on Mafalda.

Link to the web view: https://nodus.ligo.caltech.edu:30889/FE/

  8894   Mon Jul 22 21:23:04 2013 manasaUpdateGeneralVent preparation - In progress

[Gautam, Manasa]

Green steering mirrors have been swapped with PZT mirrors at the X end table. We aligned the green to the X arm.

X arm green transmission +PSL green  ~ 0.95

That's better than before the swap...woohooo

  8893   Mon Jul 22 18:49:16 2013 JenneUpdateGeneralVent preparation - In progress

I have just centered IPPOS, as well as PSL POS and PSL ANG (also called IOO POS and IOO ANG on the screens).  Annalisa is working on placing mirrors to get the IPANG beam to its QPD, so that one will be centered later.

  8892   Mon Jul 22 17:17:30 2013 annalisaUpdateendtable upgradeEnd table picture
Attachment 1: YendTable.jpg
YendTable.jpg
  8891   Mon Jul 22 17:03:25 2013 JenneUpdateGeneralMC spot positions

The results of today's MC spot position measurements:

spot positions in mm (MC1,2,3 pit MC1,2,3 yaw):
[2.3244717046516197, -0.094366247149508087, 1.6060842142158149, -0.74616561350974353, -0.67461746482832874, -1.3301448018100492]

MC1 and MC3 both have spots that are a little high in pitch, but everything else looks okay.

Actual Script:

/opt/rtcds/caltech/c1/scripts/ASS/MC/mcassMCdecenter

Plotting Script:

/opt/rtcds/caltech/c1/scripts/ASS/MC/MC_spotMeasurement_history.py

MCdecenter_22July2013.png

  8890   Mon Jul 22 16:56:14 2013 JenneUpdateGeneralVent preparation - In progress

Pre-vent checklist

  • Center all oplevs/IPPOS/IPANG
  • Align the arm cavities for IR and align Xgreen and Ygreen lasers to the arms.
    (X green+PSL green = TRX ~0.7 pre-swap and ~ 0.6 post-swap, Ygreen +PSL green TRY ~600 counts an hour or so after green was aligned to the arm.)
  • Make a record of the MC pointing
  • Align the beam at the PSL angle and position QPDs
     
  • Record good OSEM values.
     
  • Reduce input power by placing wave plate+PBS setup on the PSL table either BEFORE or AFTER THE PMC. (We will try attenuating the power using the WP + PBS that already exist after the laser. If this does not help attenuate enough, we will introduce WP+PBS after the PMC). Refer elog 6892 and elog 7299 for after the PMC detailed procedure.
     
  • Replace 10% BS before MC REFL PD with Y1 mirror and lock MC at low power.
  • Check the MC spot position measurement under the low power mode.
     
  • Close shutter of PSL-IR and green shutters at the ends
  • Make sure the jam nuts are protecting bellows
  8889   Mon Jul 22 16:30:31 2013 KojiUpdateGeneralVent preparation

[Annalisa, Manasa, Jenne, Koji]

We are working on the vent preparation.

First of all, there was no light in the interferometer.
Obviously there were lots of IFO activity in the weekend. Some were elogged, some were not.
Annalisa took her responsibility to restore the alignment and the arms recovered their flashes.

The odd thing was that the ASS got instable after we turned down the TRY PD gain from +20dB to +10dB (0dB original).
We increased the TRY gain by factor of 10 (that's the "10dB" of this PDA520. See the spec sheet) to compensate this change.
This made the ASS instable. Anyway we reduced the gain of TRY PD to 0dB. This restored the ASS.

Jenne took some more data for the QPD spectrum calibration.

Link to the vent plan

  8888   Mon Jul 22 06:58:17 2013 LisaSummaryloreAngel of the Y End Table?

Quote:

 Trying to take an image or movie of the ETMY Transmon cam, we got instead this attached image.

I think it is just some scattered green light, but others in the control room think that it is a message from somewhere or someone...

 It is not an angel, it is clearly a four leaf clover (also known as "quadrifoglio"). It is very rare, it brings good luck!

Attachment 1: image.jpg
image.jpg
  8887   Mon Jul 22 03:10:41 2013 ranaSummaryloreAngel of the Y End Table?

 Trying to take an image or movie of the ETMY Transmon cam, we got instead this attached image.

I think it is just some scattered green light, but others in the control room think that it is a message from somewhere or someone...

Attachment 1: asdasd.jpg
asdasd.jpg
  8886   Mon Jul 22 03:09:51 2013 AnnalisaUpdateGreen LockingY Arm cavity scan

Yesterday and today I was in the lab doing many cavity scan.

First I did many measurement with the cavity aligned in order to get the position of the 00 modes, then I misaligned the beam in many different ways to enhance the higher order modes.

In particular, I first misaligned the mode cleaner to make the beam clipping into the Faraday. To do this, I set to 0 the WFS gain, but I left the autolocker still enabled. In this way, the autolocker couldn't bring the mirrors back to the aligned position.

Then I misaligned also the TT2 to get even more HOMs.

Eventually, Rana came and we misaligned TT1 to clip the beam, and using TT2 we aligned back the beam to the arm.

To increase the SNR, we changed the gain of the TRY PD, setting it to 20dB (which corresponds to a factor 100 in digital scale)

I attached one scan that I did with Rana on Sunday night. I could not upload a better resolution image because the file size was too big, but here's the path to find all of the scans:

../users/annalisa/sweep/Yarm

There are many folders, one per each day I measured. In each folder there are measurements relative to aligned cavity, Pitch and Yaw misalignment.

 RXA EDIT:

The PDA520 used for TRY was set to 0 dB analog gain. This corresponds to ~500 counts out of 32768. The change to 20 dB actually increases the gain by 100. This makes the single arm lock saturate at ~25000 counts (obviously in analog before the ADC). The right setting for our usual running is probably 10 dB.

For the IMC WFS, we had disabled the turn on in the autolocker to use the IMC to steer the beam in the FI, but that was a flop (not enough range, not enough lever arm). In the end, I think we didn't get any clipping.

 

 

Attachment 1: Screenshot-scan3_0722.pdf_-_Adobe_Reader.png
Screenshot-scan3_0722.pdf_-_Adobe_Reader.png
  8885   Fri Jul 19 23:35:52 2013 KojiSummaryLSCTRY DC locking

The Y arm was locked with the TRY DC signal.

The handing off process is too complicated because there is no path from ALS to the LSC error.


 The TRY DC error signal & the gain determination

- The error signal was produced by the operation 1/SQRT(TRY) - OFFSET. The initial offset was -5.

- The sign of the TRY DC error signal depends on which side of the resonance the arm is.
  By looking at the strip chart, I determined that the sign is opposite of the ALS.
  The ALS had the gain of -25, so the TRY control gain was to be positive.

- From the strip chart on the previous entry , the slope difference between the PDH error and the TRY DC error was x500.
  The arm control with POY11 PDH had the gain of 0.2. So the target gain for the TRY DC was determined to be +100.

Handing off

- The arm was stabilized by ALS. The ALS gain was -25 with FM2/3/5/6/7/10

- YARM configuration: no trigger / no FM trigger / gain =+0 / FM5 ON / OFFSET -5

- Start handing off:
  YARM: Turned up the gain to +50

- ALS: Turned off FM6/7

- YARM: Turned on FM6/7

- ALS: Turned off FM2

- YARM: Turned on FM4

- ALS: Turned off FM3/10

- YARM: Turned on FM2/3/8/9 ON

- ALS: Reduced the gain to -15

- YARM: Increased the gain to +70

- ALS: Reduced the gain to 0

- YARM: Increased the gain to +100

HANDING OFF - DONE

Changing the offset

The offset of -5 gave the TRY of <0.1.

The detuning was reduced by giving the offset of -4. TRY went up to ~.1

The offset of -3 made TRY 0.13

The offset of -2 made TRY 0.25

The offset of -1.5 made TRY 0.4. And the arm could not be held by this error signal anymore.

 

 

Attachment 1: TRY_locking.png
TRY_locking.png
  8884   Fri Jul 19 23:35:31 2013 gautamConfigurationendtable upgradePreliminary Calibration of PZT

 I recalibrated the QPD today as I had shifted its position a little. I then identified the linear range of the QPD and performed a preliminary calibration of the Piezo tip-tilt within this range.

 


Details:

-I recalibrated the QPD as I had shifted it around a little in order to see if I could move it to a position such that I could get the full dynamic range of the piezo tilt within the linear regime of the QPD. This proved difficult because there are two reflections from the mirror (seeing as it is AR coated for 532nm and I am using a red laser). At a larger separation, these diverge and the stray spot does not bother me. But it does become a problem when I move the QPD closer to the mirror (in an effort to cut down the range in which the spot on the QPD moves). In any case, I had moved the QPD till it was practically touching the mirror, and even then, could not get the spot motion over the full range of the PZTs motion to stay within the QPD's linear regime (as verified by applying a 20Vpp 1Hz sine wave to the PZT driver board and looking at the X and Y outputs from the QPD amplifier. 

-So I reverted to a configuration in which the QPD was ~40cm away from the mirror (measured using a measuring tape).

-The new calibration constants are as follows (see attached plots):

X-Coordinate: -3.43 V/mm
Y-Coordinate: -3.41 V/mm


-I then determined the linear range of the QPD to be when the output was in the range [-0.5V 0.5V]. 

-Next, at Jenne's suggestion, I decided to do a preliminary calibration of the PZT within this linear range. I used an SR function generator to supply an input voltage to the PZT driver board's input (connected to Channel 1 of the piezo). In order to supply a DC voltage, I set a DC offset, and set the signal amplitude to 0V. I then noted the X and Y-coordinate outputs, being sure to run through the input voltages in a cyclic fashion as one would expect some hysteresis. 

-I did this for both the pitch and yaw inputs, but have only superficially analysed the latter case (I will put up results for the former later). 

 


Comments:

-There is indeed some hysteresis, though the tilt seems to vary linearly with the input voltage. I have not yet included a calibration constant as I wish to perform this calibration over the entire dynamic range of the PZT. 

 

-There is some residual coupling between the pitch and yaw motion of the tip tilt, possibly due to its imperfect orientation in the holder (I have yet to account for the QPD's tilt).

 

-I have not included a graphical representation here, but there is significantly more pitch to yaw coupling when my input signal is applied to the tip-tilts pitch input (Channel 2), as compared to when it is input to channel 1. It is not clear to me why this is so.

 

-I have to think of some smart way of calibrating the PZT over its entire range of motion, keeping the spot in the QPD's linear regime throughout. One idea is to start at one extreme (say with input voltage -10V), and then perform the calibration, re-centering the spot to 0 on the QPD each time the QPD amp output reaches the end of its linear regime. I am not sure if this will work, but it is worth a shot. The other option is to replace the red laser with a green laser (from one of the laser pointers) in the hope that multiple reflections will be avoided from the mirror. Then I will have to recalibrate the set up, and see if I can get the QPD close enough to the mirror such that the spot stays within the linear regime of the QPD. More investigation needs to be done.


Plots:

QPD Calibration Plots:
QPD-XCalib.pdf            QPD-YCalib.pdf

 

 

Piezo tilt vs input voltage plots:

 

                                                          Yaw Tilt                                                                                                                                         Pitch Tilt

Piezo_Yaw_Calib-in_QPD_linear_range-Yaw_tilt.pdf               Piezo_Yaw_Calib-in_QPD_linear_range-Pitch_tilt.pdf 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

 

  8883   Fri Jul 19 22:51:40 2013 gautamConfigurationendtable upgradeSecond mirror glued to PZT and mounted

 

I mounted the second PZT in a modified mount, and then glued a 1-inch Y2 mirror on it using superglue.

 Details:

-The mirror is a Laseroptik 1-inch, Y2 mirror with HR and AR coatings for 532 nm light.

-The procedure for mounting the mirror was the same as detailed in elog 8874. This time, I tried to orient the Piezo such that the four screws on the back face coincided with the horizontal and vertical axes, as this appeared to (somewhat) decouple the pitch and yaw motion of the tip-tilt on the first PZT.

-One thing I forgot to mention in the earlier elog: it is best to assemble the mount fully before inserting the tip-tilt into it and gluing the mirror to the tip-tilt. In particular, the stand should be screwed onto the mount before inserting the tip-tilt into the holder, as once it is in, it will block the hole through which one can screw the stand onto the mount. 

-I have placed the mirror on the SP table along with the rest of my QPD/Piezo calibration setup. I will attempt to calibrate this second PZT once I am done with the first one. 
 
Here is an image of the assembly:
Piezo-mirror_2.JPG

 

 

 

 

 

 

 

  8882   Fri Jul 19 22:35:06 2013 KojiSummaryLSCVarious Arm signal (Yarm)

The StripTool plot attached below shows various arm signals measured with the Y arm cavity swept using ALS.

Yellow: TRY

Blue: ALS additive OFFSET to the error signal

Red: Raw PDH error signal (POY11I)

Purple: Linearized PDH error (POY11/TRY)

Green: 1/Sqrt(TRY)-5 (No normalization)

Inverse Sqrt of the TRY had been implemented when this LSC controller was first coded.
It is confirmed that the calculation is working correctly.

Attachment 1: various_arm_signal.png
various_arm_signal.png
  8881   Fri Jul 19 14:04:24 2013 KojiUpdateCDSCDS FE not happy

daqd was restarted.


- tried telnet fb 8088 on rossa => same error as manasa had

- tried telnet fb 8087 on rossa => same result

- sshed into fb ssh fb

- tried to find daqpd by ps -def | grep daqd => not found

- looked at wiki https://wiki-40m.ligo.caltech.edu/New_Computer_Restart_Procedures?highlight=%28daqd%29

- the wiki page suggested the following command to run daqd /opt/rtcds/caltech/c1/target/fb/daqd -c ./daqdrc &

- ran ps -def | grep nds => already exist. Left untouched.

- Left fb.

- tried telnet fb 8087 on rossa => now it works

  8880   Fri Jul 19 12:23:34 2013 manasaUpdateCDSCDS FE not happy

I found CDS rt processes in red. I did 'mxstreamrestart' from the medm. It did not help. Also ssh'd into c1iscex and tried 'mxstreamrestart' from the command line. It did not work either.

I thought restarting frame builder would help. I ssh'd to fb. But when I try to restart fb I get the following error:

controls@fb ~ 0$ telnet fb 8088
Trying 192.168.113.202...
telnet: connect to address
192.168.113.202: Connection refused

 

Screenshot-Untitled_Window.png

  8879   Fri Jul 19 12:02:18 2013 manasaUpdateLSCPRMI+Y arm ALS success!

Data retrieved using getdata (30 minutes trend) saved at

/users/manasa/data/130717/PRMI_YALS

  8878   Fri Jul 19 12:00:12 2013 manasaUpdateGreen LockingALS Y performance with the new whitening filter

Quote:


(Manasa downloaded the 2k sampled data so that we can use this for presentations.)

 Path to data (retreived using getdata)

/users/manasa/data/130717/YALS_scan

  8877   Thu Jul 18 23:34:40 2013 gautamConfigurationendtable upgradeCoarse adjustment of PZT axes orientation in mount

 I have managed to orient the PZT in the mount such that its axes are approximately aligned with the vertical and the horizontal. 

In the process, I discovered that the 4 screws on the back face of the PZT correspond to the location of the piezoelectric stacks beneath the tip-tilt platform. The PZT can therefore be oriented during the mounting process itself, before the mirror is glued onto the tip-tilt platform.

In order to verify that the pitch and yaw motion of the mirror have indeed been roughly decoupled, I centred the spot on the QPD, fed to the 'pitch' input of the PZT driver board (connected to channel 1 of the PZT) a 10 Vpp, 1 Hz sine wave from the SR function generator (having turned all the other relevant electronics, HV power supply etc ON. The oscilloscope trace of the output observed on the QPD is shown. The residual fluctuation in the Y-coordinate (blue trace) is I believe due to the tilt in the QPD, and also due to the fact that the PZT isnt perfectly oriented in the mount.

It looks like moving the tip-tilt through its full range of motion takes us outside the linear regime of the QPD calibration. I may have to rethink the calibration setup to keep the spot on the QPD in the linear range if the full range is to be calibrated, possibly decrease the distance between the mirror and the QPD. Also, in the current orientation, CH1 on the PZT controls YAW motion, while CH2 controls pitch.

Oscilloscope Trace:

Yellow: X-coordinate

Blue: Y-coordinate

PZT_out_waveform.jpg

  8876   Thu Jul 18 21:45:36 2013 CharlesUpdateISSISS - Full Schematic

 Here I have included the full schematic (so far) of the proposed ISS. There are two sheets: the first schematic details the filter stages and their accompanying circuitry while the second schematic details the RMS threshold detection and subsequent triggering.

The first schematic is fairly self explanatory as to what different portions do, and I have annotated much of the second schematic as there are some non-traditional components etc.

I have not yet included some mechanism to adjust the threshold voltage in real time or any of the power regulation, but these should follow fairly quickly.

Attachment 1: 40mServo_v1.pdf
40mServo_v1.pdf 40mServo_v1.pdf
  8875   Thu Jul 18 21:12:58 2013 gautamConfigurationendtable upgradeAI Board-D000186-All channels modified

 

 I carried some further modifications and tests to the AI Board. Details and observations here:

  • I switched out the resistors for all the remaining 7 channels, using the same substitutions as detailed here
  • I then verified that the modified transfer function for all 8 channels using the SR785. I did not collect data for all the channels as netgpib was taking ages, but I did use the cursor on the screen to verify the position of the first notch at ~64 kHz. I noticed that all the channels did not have the lowest point of the notch at the same frequency. Rather, (at least on the screen), this varied between 63kHz and 67kHz. I would put this down to component tolerance. Assuming 5% tolerance shifts the theoretical notch frequency from 66268 Hz to 63112 Hz. 
  • After verifying the transfer functions, I went to 1Y4 and plugged the AI board into the eurocrate. I then connected the input of the AI board to the DAC output using my custom ribbon cable. Next, I used the excitation points set up earlier to send a 1 kHz, 32000 counts amplitude sine wave through the channels one at a time. I monitored the output using an oscilloscope and the LEMO monitor channels on the front panel of the board.
  • I found that the single-ended output of the AI board swings between -10 V and 10 V (w.r.t ground, oscilloscope trace attached). This is good because this is the range of input voltage to the PZT driver boards required to realize the full actuation range of the PZTs.
  • I also verified that the connections on the custom ribbon cable are correct (channel map was right) and that there were no accidental shorts (I checked other channels' output monitor while driving one channel). 

I think the board is okay to be used now.

AI_Board_Output.jpg

 

  8874   Thu Jul 18 20:20:52 2013 gautamConfigurationendtable upgradeFirst mirror glued to PZT and mounted in modified mounts

 

 Yesterday, I mounted the first PZT in one of the modified mounts, and then glued a 2-inch Y2 mirror on it using superglue.

Details:

-The mirror is a 2-inch, Y2 mirror with HR and AR coatings for 532 nm light.

-The AR side of the mirror had someone's fingerprint on it, which I removed (under Manasa's guidance) using tweezers wrapped in lens cleaning paper, and methanol.

-Before gluing the mirror, I had to assemble the modified mount. Manasa handed over the remaining parts of the mounts (which are now in my newly acquired tupperware box along with all the other Piezo-related hardware). I took the one labelled A, and assembled the holder part. I then used one of the new mounts (2.5 inches, these are with the clean mounts in a cardboard box in the cupboard holding the green optics along the Y-arm) and mounted the holder on it. 

-Having assembled the mount, I inserted the piezo tip-tilt into the holder. The wedge that the machine shop supplied is useful (indeed required) for this. 

-I then cleaned the AR surface of the mirror and the top-surface of the tip-tilt. 

-The gluing was done using superglue which Steve got from the bookstore (the remaining tube is in the small fridge). We may glue the other mirror using epoxy. I placed 4 small drops of superglue on the tip-tilt's top surface, placed the mirror with its AR face in contact with the piezo, and applied some pressure for a short while until the glue spread out fairly evenly. I then left the whole setup to dry for about half an hour.

-Steve suggested using a reference piece (I used two small bolts) to verify when the glue had dried.

-Finally, I attached the whole assembly to a base.

Here it is in action in my calibration setup (note that it has not been oriented yet. i.e. the two perpendicular axes of the piezo are for the time being arbitrarily oriented. And maybe the spreading of the glue wasn't that even after all...):

Piezo-mirror.jpg

 

Sidenote:

Yesterday, while setting stuff up, I tested the piezo with a 0.05 Hz, 10Vpp input from the SR function generator just to see if it works, and also to verify that I had set up all my electronics correctly. Though the QPD was at this point calibrated, I did observe periodic motion of both the X and Y outputs of my QPD amp! Next step- calibration... 

  8873   Thu Jul 18 19:09:08 2013 gautamConfigurationendtable upgradeQPD Calibration for PZT Calibration

Summary 

I have been working on setting up a QPD which can eventually be used to calibrate the PZT, and also orient the PZT in the mount such that the pitch and yaw axes roughly coincide with the vertical and horizontal.

The calibration constants have been determined to be:

X-axis: -3.69 V/mm

Y-axis: -3.70V/mm

Methodology:

I initially tried using the QPD setup left behind by Chloe near MC2, but this turned out to be dysfunctional. On opening out the QPD, I found that the internal circuitry had some issues (shorts in the wrong places etc.) Fortunately, Steve was able to hand me another working unit. For future reference, there are a bunch of old QPDs which I assume are functional in the cabinet marked 'Old PDs' along the Y-arm. 

I then made a circuit with which to read out the X and Y coordinates from the QPD. This consists of 4 buffer amplifiers (one for each quadrant), and 3 summing amplifiers (outputs are A+B+C+D = sum, B+C-A-D = Y-coordinate, and A+B-C-D = X-coordinate) that take the appropriate linear combinations of the 4 quadrants to output a voltage that may be calibrated against displacement of the QPD. 

The output from the QPD is via a sub-D connector on the side of the pomona box enclosing the PD and the circuitry, with 7 pins- 3 for power lines, and 4 for the 4 quadrants of the QPD. It was a little tricky to figure the pin-out for this connector, as there was no way to use continuity checking to map the pins to quadrants. Therefore, I used a laser pointer, and some trial and error (i.e. shine the light on a given quadrant, and check the sign of the X and Y voltages on an oscilloscope) to map the pin outs. Steve tells me that these QPDs were made long before colour code standardisation, but I note here the pin outs in any case for future reference (the quadrant orientations are w.r.t the QPD held with all the circuitry above it, with the active surface facing me):

Red= +Vcc

Black= -Vcc

Green = GND

Blue = Upper Left Quadrant

White = Upper Right Quadrant

Purple = Lower Left Quadrant

Grey = Lower Right Quadrant

Chloe had noted that there was some issue with the voltage regulators on her circuit (overheating) but I suspect this may have been due to the faulty internal circuitry. Also, she had used 12 V regulators. I checked the datasheet of the QPD, Op-Amp LF347 (inside the pomona box) and the OP27s on my circuit, and found that they all had absolute maximum ratings above 18V, so I used 15V voltage regulators. The overheating problem was not a problem anymore.

I then proceeded to arrange a set up for the calibration (initially on the optical bench next to MC2, but now relocated to the SP table, and a cart adjacent to it). It consists of the following:

  • He-Ne laser source
  • Y2 2-inch mirror (AR and HR coated for 532nm) glued onto the PZT and mounted on a machined Newport U100P  - see this elog for details.
  • QPD mounted on a translational stage whose micrometers are calibrated in tenths of an inch (in the plots I have scaled this to mm)
  • A neutral density filter (ND = 2.0) which I added so that the QPD amplifier output did not saturate. I considered using a lens as well to reduce the spot size on the QPD but found that after adding the ND filter, it was reasonably small.
  • High-voltage power supply (on cart)
  • Two SR power supplies (for the PZT driver board and my QPD amplifier
  • SR function generator
  • Laser power source
  • Two oscilloscopes
  • Breadboard holding my QPD amplifier circuit

Having set everything up and having done the coarse alignment using the mirror mount, I proceeded to calibrate the X and Y axes of the QPD using the translational stage. The steps I followed were:

  • Centre spot on QPD using coarse adjustment on the mirror mount: I gauged this by monitoring the X and Y voltage outputs on an oscilloscope, and adjusted things till both these went to zero.
  • Used the tilt knob on the translational stage to roughly decouple the X and Y motion of the QPD.
  • Kept Y-coordinate fixed, took the X-coordinate to close to its maximum value (I gauged this by checking where the voltage stopped changing appreciably for changes in the QPD position.
  • Using this as a starting point, I moved the QPD through its X range, noting voltage output of the X-coordinate (and also the Y) on an oscilloscope.
  • Repeated the procedure for the Y-coordinate.
  • Analysis follows largely what was done in these elogs. I am attaching the script I used to fit an error function to the datapoints, this is something MATLAB should seriously include in cftool (note that it is VERY sensitive to the initial guess. I had to do quite a bit of guessing).

The plots are attached, from which the calibration values cited above are deduced. The linear fits for the orthogonal axis were done using cftool. There is some residual coupling between the X and Y motions of the QPD, but I think this os okay my purposes. 

My next step would be to first tweak the orientation of the PZT in the mount while applying a small excitation to it in order to decouple the pitch and yaw motion as best as possible. Once this is done, I can go ahead and calibrate the angular motion of the PZT in mrad/V.

 

                                                            X-Axis                                                                                                                                Y-axis                                                

QPD-XCalib.pdf  QPD-YCalib.pdf

 

 

Attachment 3: Error_Function_Fitting.zip
  8872   Thu Jul 18 16:30:08 2013 KojiUpdateLSCPRMI+Y arm ALS Sensing Matrices

Hmm. I agree that something was funny.
Let's take the matrix without the arms and confirm the measurement is correct.

ELOG V3.1.3-